Show work, thanks

2.5.1. The probability function of a random variable Y is given by p (i)-% 1 0, 1, 2, , where λ is a known positive value and c is a constant. (a) Find c. (b) Find P(Y 0) (c) Find P(Y > 2).
Answer #1

(a)

c is got by noting that the Total Probability = 1

We get:

\sum_{i=0}^{\infty }\frac{c\lambda ^{i}}{i!}=1

i.e.,

c\sum_{i=0}^{\infty }\frac{\lambda ^{i}}{i!}=1 (1)

By Power Series, we have:

\sum_{i=0}^{\infty }\frac{\lambda ^{i}}{i!}=e^{\lambda } (2)

Substituting (2), equation (1) becomes:

c\times e^{\lambda }=1

So,

c= e^{-\lambda }

So,

Answer is:

e^{-\lambda }

(b)

So, the probability function of Y is given by:

p(i)=\frac{e^{-\lambda }\lambda ^{i}}{i!}

P(Y=0)=\frac{e^{-\lambda }\lambda ^{0}}{0!}=e^{-\lambda }

So,

Answer is:

P(Y=0)=e^{-\lambda }

(c)
P(Y>2) = 1 -[P(Y=0)+P(Y=1)+P(Y=2)] (3)

P(Y=0)=e^{-\lambda } (4)

P(Y=1)=\frac{e^{-\lambda }\lambda ^{1}}{1!}=\lambda e^{-\lambda } (5)

P(Y=2)=\frac{e^{-\lambda }\lambda ^{2}}{2!}=\frac{e^{-\lambda }\lambda ^{2}}{2} (6)

Substituting (4), (5) and (6), equation (3) becomes:

P(Y>2)=1-[e^{-\lambda }+\lambda e^{-\lambda }+\frac{e^{-\lambda }\lambda ^{2}}{2}]

Thus, we get the answer as:

P(Y>2)=1-e^{-\lambda }[1+\lambda +\frac{\lambda ^{2}}{2}]